# On Scaling Up 3D Gaussian Splatting Training

### v1

@@ Appendix C Additional Experiments Setting and Statistics
C.1 Statistics for Mip-NeRF 360, Tank&Temples dataset and DeepBlending
@@
C.2 Scalability
Table 3, 4 and 5 show the increased reconstruction quality with more gaussians. While many hyperparameters influence the number of Gaussians created by densification, we focused on adjusting three key parameters: (1) the stop iteration for densification, (2) the threshold for initiating densification, and (3) the threshold for deciding whether to split or clone a Gaussian. Initially, we gradually increased the densification stop iteration to 5,000 iterations. However, due to the pruning mechanism, this adjustment alone proved insufficient. Consequently, we also lowered the two thresholds to generate more Gaussians. For a fair comparison, all other densification parameters—such as the interval, start iteration, and opacity reset interval—were kept constant. For the Rubble scene, each experiment run for the same 125 epochs, exposing models to 200,000 images, ensuring consistency. Although training larger models for longer durations and lowering the positional learning rate improved results in my observations, we maintained consistent training steps and learning rates across all experiments to ensure fairness.
Table 6, 7 show the Throughput Scalability by Increasing batch size and leveraging more GPUs, for Rubble and Train scene, respectively. Essentially, more GPUs and larger batch size give higher throughput.
Table 8 demonstrates that additional GPUs increase available memory for more Gaussians, evaluated on the Rubble scene with various batch sizes reflecting different levels of activation memory usage. Essentially, more GPUs provide additional memory to store Gaussians, while a larger batch size increases activation memory usage, leaving less memory available for Gaussians."

### v2

@@ B.2 Gaussian Distribution Rebalance
B.3 Scaling Rule Hyperparameter Ablation
The effectiveness of our automatic hyperparameter scaling rules is demonstrated in the ablation study, as shown in Figure 12.
(a) Learning rate scaling rules vs. BS invariance.
(b) Momentum scaling rules vs. BS invariance.
Figure 12: We plot the training trajectories of the diffuse color parameters on “Rubble”, when training with batch size $\in[4,16,32]$ using different learning rate and momentum scaling strategies. Cumulative weight updates using the square-root learning rate scaling rule (a, red curves) and exponential momentum scaling rule (b, red curves) maintain high cosine similarity to batch-size 1 updates and have norms that are roughly invariant to the batch size. All trajectories in (a) employ our proposed exponential momentum scaling but differing learning rate scaling; while all trajectories in (b) employ our proposed square-root learning rate scaling but differing momentum scaling.
@@ Appendix C Additional Experiments Setting and Statistics
C.1 Ablation Study
Figure 10 illustrates that our load balancing techniques and increased batch size significantly improve training throughput on 1080p Mip-NeRF360 dataset, compared to the one GPU baseline and our straightforward distributed system with a conventional batch size of one and no load balancing. Similar results are observed with the 4K Rubble Dataset, as shown in Figure 15. Although good speed can be achieved without load balancing, load balancing allows us to consistently achieve even higher throughput across various types and scales of scenes. The ablation study for the learning rate scaling strategies have already been discussed in 4.2, along with our analysis.
C.2 Statistics for Mip-NeRF 360, Tank&Temples and DeepBlending datasets
We provide full statistics of training results on Mip-NeRF 360, Tank&Temples and DeepBlending datasets in Table 3.
@@ C.2 Statistics for Mip-NeRF 360, Tank&Temples and DeepBlending datasets
C.3 Scalability
Table 4, 5 and 6 show the increased reconstruction quality with more gaussians. While many hyperparameters influence the number of Gaussians created by densification, we focused on adjusting three key parameters: (1) the stop iteration for densification, (2) the threshold for initiating densification, and (3) the threshold for deciding whether to split or clone a Gaussian. Initially, we gradually increased the densification stop iteration to 5,000 iterations. However, due to the pruning mechanism, this adjustment alone proved insufficient. Consequently, we also lowered the two thresholds to generate more Gaussians. For a fair comparison, all other densification parameters—such as the interval, start iteration, and opacity reset interval—were kept constant. For the Rubble scene, each experiment run for the same 125 epochs, exposing models to 200,000 images, ensuring consistency. Although training larger models for longer durations and lowering the positional learning rate improved results in my observations, we maintained consistent training steps and learning rates across all experiments to ensure fairness.
Table 7, 8 show the Throughput Scalability by Increasing batch size and leveraging more GPUs, for Rubble and Train scene, respectively. Essentially, more GPUs and larger batch size give higher throughput. More GPUs provide more computational power while larger batch size can utilze these GPUs better.
Table 9 demonstrates that additional GPUs increase available memory for more Gaussians, evaluated on the Rubble scene with various batch sizes reflecting different levels of activation memory usage. We can achieve linear scaling as illustrated in Figure 15. Essentially, more GPUs provide additional memory to store Gaussians, while a larger batch size increases activation memory usage, leaving less memory available for Gaussians.
@@ 2 GPU OOM
Method Rubble Building MatrixCity Block_All
Total Time Time Decomposition Total Time Time Decomposition Total Time Time Decomposition
CityGS Official 2.88 hrs train_coarse: 43min14s 4.57 hrs train_coarse: 44min3s 8.25 hrs train_coarse: 47min31s
data_partition: 6min5s data_partition: 17min37s data_partition: 1h28min22s
9 cells train on 4 A100: 2h3min 20 cells train on 4 A100: 3h31min 36 cells train on 4 A100: 5h57min
Merge point cloud: 59s Merge point cloud: 1min21s Merge point cloud: 2min27s
CityGS 200K-images 2.18 hrs train_coarse: 43min14s 2.22 hrs train_coarse: 44min3s 3.60 hrs train_coarse: 47min31s
data_partition: 6min5s data_partition: 17min37s data_partition: 1h28min22s
9 cells train on 4 A100: 1h20min 20 cells train on 4 A100: 1h11min 36 cells train on 4 A100: 1h20min
Merge point cloud: 57s Merge point cloud: 1min24s Merge point cloud: 1min9s
Grendel 200K-images (ours) 0.85 hrs trained on 4 A100: 51min 0.90 hrs trained on 4 A100: 54min 1.22 hrs trained on 4 A100: 1h13min
Table 10: Time Decomposition for experiments in Table 2. CityGaussian employs a divide-and-conquer approach with four steps: coarse training, partitioning the scene into cells, training each cell independently, and merging the resulting point clouds. In contrast, 3DGS on multiple GPUs with Grendel can be run in the same way as the original single-GPU training, simply by allocating additional GPUs.
C.4 Render Speedup
We compare Grendel’s rendering speeds on single-GPU and 4-GPU setups for the Rubble, MatrixCity Block-All, and Bicycle scenes, using our state-of-the-art trained Gaussian models corresponding to Figure 11. We experiment with various resolutions and scene scales, rendering one image at a time for fair comparison. The 4-GPU setup achieves a speedup of 1.88x to 2.63x, as shown in Table 11.
Scene (Resolution) 1 GPU 4 GPU
Rubble (3436x4591) 8.8 img/s 21.7 img/s
Matrixcity Block-all (1080x1920) 29.2 img/s 76.8 img/s
Bicycle (1275x1920) 42.6 img/s 80.3 img/s
Table 11: 4-GPU Render Speedup with Grendel Compared to Single-GPU Rendering
Appendix D Additional Related Works
Distributed training for neural networks. Existing work have exploited various types of parallelism to train neural networks across GPUs. These include data parallelism (Li et al., 2020), tensor parallelism (Shoeybi et al., 2020; Narayanan et al., 2021), pipeline parallelism (Huang et al., 2019; Narayanan et al., 2019) and FSDP (Zhao et al., 2023; Rajbhandari et al., 2020). Several systems also support multiple types parallelism and/or aim to automatically partition the workload to optimize speed (Zheng et al., 2022; Xu et al., 2021; Wang et al., 2019). However, as we discussed earlier (§1), neural network and 3DGS have very different computation patterns. The former performs repeated layer-wise computation dominated by dense matrix multiply operations while the latter’s 3 stages training process is irregular and sparse. As a result, although Grendel’s distribution strategy may resemble those seen in existing work (e.g., FSDP Zhao et al. (2023)), the details are quite different.